### v1

@@ 4.4. Higher $\mathrm{L}^{p}$-based Sobolev regularity
Having established the well-posedness of (4.8), we are now ready to derive higher-regularity results directly for any $p>1$. The proof strategy is a generalization of the trick used in [61].
@@
Let $1<p<\infty$, and assume $\Gamma$ is of class $C^{m+3}$ for some nonnegative integer $m$. If $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$, then the solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of (4.8) (provided by 4.8) belongs to $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Moreover, there exists a positive constant $C$, depending only on $\Gamma$, $m$ and $p$ such that
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq C\left(\|\mathbf{f}\|_{m,p,\Gamma}+\|g\|_{m+1,p;\Gamma}\right).$ (4.20)
We first proof the case $m=0$. For general $m$ the proof will follow by induction.
@@
Step 1: Case $m=0$. Recalling 2.1 (and using integration by parts and density arguments), we start by choosing $\mathbf{v}=\nabla_{\Gamma}q$ with arbitrary $q\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ in the first equation of (4.8), thereby arriving at
@@
for each $q\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$. Furthermore, upon using the fact that $\operatorname{div}_{\Gamma}\mathbf{u}=g$ a.e. on $\Gamma$, and integrating-by-parts the last term in the right-hand side, we further obtain
$\int_{\Gamma}\uppi\,\Delta_{\Gamma}q=\int_{\Gamma}\mathbf{h}[\mathbf{u}]\cdot\nabla_{\Gamma}q,\qquad\forall q\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma),$ (4.21)
where $\mathbf{h}[\mathbf{u}]:=\mathbf{f}+\nabla_{\Gamma}g+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$. That is, $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ satisfies the ultra-weak formulation of the Laplace–Beltrami (3.11) with right-hand side given by $\operatorname{div}_{\Gamma}\mathbf{h}[\mathbf{u}]$, understood as an element of $(\mathrm{W}^{2,p^{*}}_{\#}(\Gamma))^{\prime}$ by integration by parts. Since $\operatorname{div}_{\Gamma}\mathbf{h}[\mathbf{u}]$ can also be seen as an element of $(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$, we deduce from 3.16 (well-posedness of the Laplace–Beltrami operator in $\mathrm{W}^{1,p}_{\#}(\Gamma)$) that $\uppi\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ and
@@
With this further regularity on $\uppi$, we can now integrate by parts in the first equation of (4.8), thus arriving at
@@
where $\mathbf{j}[\uppi]:=\mathbf{f}-\nabla_{\Gamma}\uppi\in\mathbf{L}^{p}_{t}(\Gamma)$. Therefore, by the $\mathrm{L}^{p}$-based regularity of the Bochner–Laplace operator (see Section 3.4), we infer that $\mathbf{u}\in\mathbf{W}^{2,p}_{t}(\Gamma)$ and
@@
In this way, combining (4.22) and (4.23), and using 4.8 (well-posedness of tangent Stokes in $\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$) to bound $\|\mathbf{u}\|_{0,p;\Gamma}$ in terms of $\mathbf{f}$ and $g$, we finally arrive at
@@
Step 2: Case $m\geq 1$. Assume the Theorem holds for a fixed $m\geq 0$. Assume $\Gamma$ is of class $C^{m+4}$ and $\mathbf{f}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$. By the induction hypothesis we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq C_{m,p}\left(\|\mathbf{f}\|_{m,p,\Gamma}+\|g\|_{m+1,p;\Gamma}\right).$ (4.24)
@@
Now, proceeding as in the proof of the case $m=0$, we observe that $\mathbf{h}[\mathbf{u}]:=\mathbf{f}+\nabla_{\Gamma}g+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{u}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$. Hence, by the $\mathrm{L}^{p}$-based Sobolev regularity of the Laplace–Beltrami operator (see Section 3), we deduce that $\uppi\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ and
@@
In turn, $\mathbf{j}[\uppi]:=\mathbf{f}-\nabla_{\Gamma}\uppi\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$, whence by the $\mathrm{L}^{p}$-based regularity of the Bochner–Laplace operator (see Section 3.4), we obtain that $\mathbf{u}\in\mathbf{W}^{m+3,p}_{t}(\Gamma)$ and
This concludes the induction step, and hence, the proof of 4.9. ∎
5. The Stokes operator
In this section we study spectral properties of the tangent Stokes operator on $\Gamma$ induced by (S), thereby mimiching spectral properties of the classical Stokes operator in flat domains [19, §4] [63, Part I, §2]. We will later make use of this theory in Section 6.2 to establish the existence of solutions for the incompressible tangent Navier–Stokes equations via the Faedo–Galerkin method.
Definition 5.1 (Stokes operator on $\Gamma$).
Let $1<p<\infty$ and assume $\Gamma$ is of class $C^{3}$. We define the Stokes operator on $\Gamma$ as the operator $A:\mathbf{W}^{2,p}_{t,\sigma}(\Gamma)\rightarrow\mathbf{L}^{p}_{t,\sigma}(\Gamma)$ by $A\mathbf{v}=-P_{\sigma}(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{v}$). For $p=2$ we write $D(A):=\mathbf{H}^{2}_{t,\sigma}(\Gamma)$.
By properties of the Leray projection (cf. 4.4), it is immediate to see that $\|A\mathbf{v}\|_{0,p;\Gamma}\leq C_{p}\|\mathbf{v}\|_{2,p;\Gamma}$ for each $\mathbf{v}\in\mathbf{W}^{2,p}_{t,\sigma}(\Gamma)$, where $C_{p}>0$ is a constant depending on $\Gamma$ and $p$. If $\Gamma$ is of class $C^{m+3}$ for some $m\in\mathrm{N}\cup\{0\}$, 4.3 implies that the operator $A$ maps $\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$ to $\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$. Furthermore, for each $\mathbf{u}\in\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$ and $\mathbf{f}\in\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$, it follows from the definition of the Stokes operator that $A\mathbf{u}=\mathbf{f}$ if and only if there exists $\uppi\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ such that
@@ 5. The Stokes operator
Then, by 4.9 ($\mathrm{L}^{p}$-based Sobolev regularity of the tangent Stokes equations) we deduce that $A$ is a bijection from $\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$ to $\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$ and
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\lesssim\|\mathbf{f}\|_{m,p,\Gamma}=\|A\mathbf{u}\|_{m,p,\Gamma}.$ (5.1)
Moreover, the Bounded Inverse Theorem [12, Corollary 2.7], [31, Teorema 3.4] implies that
$\|A\mathbf{u}\|_{m,p,\Gamma}\lesssim\|\mathbf{u}\|_{m+2,p;\Gamma}.$ (5.2)
@@
defines an equivalent inner product in $\mathbf{H}^{m+2}_{t,\sigma}(\Gamma)$. On the other hand, if $m+2=2\ell+1$ is odd we also use 3.8 (Poincaré inequality) to deduce that
@@
defines an equivalent inner product in $\mathbf{H}^{m+2}_{t,\sigma}(\Gamma)$.
Proposition 5.2 (properties of the Stokes operator).
The Stokes operator $A$ is symmetric and positive with respect to the $\mathbf{L}^{2}(\Gamma)$-inner product. More precisely, for each $\mathbf{u}\in\mathbf{H}^{2}_{t,\sigma}(\Gamma)$ and $\mathbf{v}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$,
$\int_{\Gamma}A\mathbf{u}\cdot\mathbf{v}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v},\qquad\text{and}\qquad\int_{\Gamma}A\mathbf{u}\cdot\mathbf{u}\geq 0.$ (5.5)
Proof.
Since $P_{\sigma}\mathbf{u}=\mathbf{u}=\mathbf{P}\mathbf{u}$ and $P_{\sigma}\mathbf{v}=\mathbf{v}=\mathbf{P}\mathbf{v}$, we have by the integration-by-parts formula [11, Lemma B.5] (valid by density) that
$\int_{\Gamma}A\mathbf{u}\cdot\mathbf{v}=-\int_{\Gamma}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{u}\cdot\mathbf{v}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v},$
whose right-hand side is symmetric in $\mathbf{u}$ and $\mathbf{v}$. The positivity of $A$ follows from taking $\mathbf{v}=\mathbf{u}$. ∎
Theorem 5.3 (eigendecomposition of the Stokes operator).
Assume $\Gamma$ is of class $C^{m+3}$ ($m\geq 0$). Then, there exists a family of functions $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}\subseteq\bigcap_{2\leq p<\infty}\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$ and a nondecreasing sequence of $\{\omega_{n}\}_{n\in\mathrm{N}}$ of positive numbers with $\lim_{n\to\infty}\omega_{n}=\infty$ such that

### v2

@@ 4.3. Higher $\mathrm{L}^{p}$-based Sobolev regularity
Having established the well-posedness of (4.1), we are now ready to derive higher-regularity results directly for any $p\in(1,\infty)$. The proof strategy is a generalization of the trick used in [53].
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m$. If $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$, then the solution $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ of (4.1) (provided by 4.7) belongs to $\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Moreover, there exists a positive constant $C$, depending only on $\Gamma$, $m$ and $p$, such that
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq C\left(\|\mathbf{f}\|_{m,p;\Gamma}+\|g\|_{m+1,p;\Gamma}\right).$ (4.14)
@@
We first prove the case $m=0$. For general $m$ the proof will follow by induction.
Step 1: Case $m=0$. Recalling 2.13 (weak commutator identity), and using a density argument, we start by choosing $\mathbf{v}=\nabla_{\Gamma}q$ with arbitrary $q\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ in the first equation of (4.1), thereby arriving at
@@
Furthermore, upon using the fact that $\operatorname{div}_{\Gamma}\mathbf{u}=g$ a.e. on $\Gamma$, and integrating-by-parts (cf. (2.5)) the second term in the right-hand side, we further obtain
$\int_{\Gamma}\uppi\,\Delta_{\Gamma}q=\int_{\Gamma}\mathbf{h}[\mathbf{u}]\cdot\nabla_{\Gamma}q,\qquad\forall q\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma),$
where $\mathbf{h}[\mathbf{u}]:=\mathbf{f}+\nabla_{\Gamma}g+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{u}\in\mathbf{L}^{p}_{t}(\Gamma)$. That is, $\uppi\in\mathrm{L}^{p}_{\#}(\Gamma)$ satisfies the ultra-weak formulation of the Laplace–Beltrami (2.10) with right-hand side given by $\operatorname{div}_{\Gamma}\mathbf{h}[\mathbf{u}]$, understood as an element of $(\mathrm{W}^{2,p^{*}}_{\#}(\Gamma))^{\prime}$ by integration by parts. Since $\operatorname{div}_{\Gamma}\mathbf{h}[\mathbf{u}]$ can also be seen as an element of $(\mathrm{W}^{1,p^{*}}_{\#}(\Gamma))^{\prime}$, we deduce from 2.15 (Laplace–Beltrami) that $\uppi\in\mathrm{W}^{1,p}_{\#}(\Gamma)$ and
@@
With this further regularity on $\uppi$, we can now integrate by parts in the first equation of (4.1), yielding
@@
where $\mathbf{j}[\uppi]:=\mathbf{f}-\nabla_{\Gamma}\uppi\in\mathbf{L}^{p}_{t}(\Gamma)$. Therefore, by 3.7 ($\mathbf{W}^{m+2,p}$-regularity of the Bochner Laplacian on $\Gamma$), we infer that $\mathbf{u}\in\mathbf{W}^{2,p}_{t}(\Gamma)$ and
@@
In this way, combining (4.15) and (4.16), and using 4.7 (well-posedness of tangent Stokes in $\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$) to bound $\|\mathbf{u}\|_{0,p;\Gamma}$ in terms of $\mathbf{f}$ and $g$, we finally arrive at
@@
Step 2: Case $m\geq 1$. Assume the assertion holds for a fixed $m\geq 0$. Assume $\Gamma$ is of class $C^{m+3,1}$ and $\mathbf{f}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$ and $g\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$. By the induction hypothesis we deduce that $(\mathbf{u},\uppi)\in\mathbf{W}^{m+2,p}_{t}(\Gamma)\times\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\leq C_{m,p}\left(\|\mathbf{f}\|_{m,p;\Gamma}+\|g\|_{m+1,p;\Gamma}\right).$ (4.17)
@@
Now, proceeding as in the proof of the case $m=0$, we observe that $\mathbf{h}[\mathbf{u}]:=\mathbf{f}+\nabla_{\Gamma}g+(\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{u}\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$. Hence, by the higher regularity result within 2.15 (Laplace–Beltrami) and using (4.17) we deduce that $\uppi\in\mathrm{W}^{m+2,p}_{\#}(\Gamma)$ and
$\|\uppi\|_{m+2,p;\Gamma}\lesssim\|\operatorname{div}_{\Gamma}\mathbf{h}[\mathbf{u}]\|_{m,p;\Gamma}\lesssim\|\mathbf{f}\|_{m+1,p;\Gamma}+\|g\|_{m+2,p;\Gamma}+\|\mathbf{u}\|_{m+1,p;\Gamma}\lesssim\|\mathbf{f}\|_{m+1,p;\Gamma}+\|g\|_{m+2,p;\Gamma}.$
In turn, $\mathbf{j}[\uppi]:=\mathbf{f}-\nabla_{\Gamma}\uppi\in\mathbf{W}^{m+1,p}_{t}(\Gamma)$, whence by 3.7 ($\mathbf{W}^{m+2,p}$-regularity of the Bochner Laplacian on $\Gamma$) we obtain that $\mathbf{u}\in\mathbf{W}^{m+3,p}_{t}(\Gamma)$ and
This concludes the induction step, and hence, the proof of 4.8. ∎
4.4. The Stokes operator
In this section, we study spectral properties of the tangent Stokes operator on $\Gamma$ induced by (S), thereby mimicking spectral properties of the classical Stokes operator in flat domains [18, §4] [55, Part I, §2]. We will later make use of this theory in Section 5.2 to establish the existence of solutions for the incompressible tangent Navier–Stokes equations via the Faedo–Galerkin method. We start by introducing the Stokes operator under the “minimal regularity” regime.
Definition 4.9 (Stokes operator on $\Gamma$).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. We define the Stokes operator on $\Gamma$ as the linear operator $\widetilde{A}:\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ such that
$\langle\widetilde{A}\mathbf{u},\mathbf{v}\rangle_{(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}\times\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)}:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v},\qquad\forall\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma),\quad\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$ (4.18)
In agreement with our philosophy of requiring minimal regularity of $\Gamma$, we favor 4.9 in lieu of the classical approach of defining a “strong” Stokes operator $A=P_{\sigma}(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\cdot)$ because the latter would require $\Gamma$ to be of class $C^{2,1}$. We reveal the connection between $\widetilde{A}$ and $A$ in (4.24) below.
From 4.5 (injectivity of the adjoint operator), 4.6 (surjectivity of the adjoint operator) and [5, Theorem A.2] (Banach–Nečas–Babuška) we know that $\widetilde{A}:\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ is a continuous bijection. For $p=2$, we notice that endowing the Hilbert space $H:=\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ and its dual $H^{\prime}$ with the equivalent inner products (cf. 3.3 (Poincaré inequality))
$(\mathbf{u},\mathbf{v})_{H}:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v},\qquad\forall\mathbf{u},\mathbf{v}\in H.$ (4.19)
and
$(\mathbf{f},\mathbf{g})_{H^{\prime}}:=(\widetilde{A}^{-1}\mathbf{f},\widetilde{A}^{-1}\mathbf{g})_{H}=\langle\mathbf{f},\widetilde{A}^{-1}\mathbf{g}\rangle_{H^{\prime}\times H}=\langle\mathbf{g},\widetilde{A}^{-1}\mathbf{g}\rangle_{H\times H^{\prime}},\qquad\forall\,\mathbf{f},\mathbf{g}\in H^{\prime},$ (4.20)
trivially turns $\widetilde{A}:H\rightarrow H^{\prime}$ into an isometry. Let $\iota:H\rightarrow H^{\prime}$ be the compact canonical embedding of $H$ into $H^{\prime}$ given by (cf. Section 2.1)
$\langle\iota(\mathbf{u}),\mathbf{v}\rangle_{H^{\prime}\times H}:=\int_{\Gamma}\mathbf{u}\cdot\mathbf{v},\qquad\forall\mathbf{u},\mathbf{v}\in H.$
We next introduce the linear operator $T:H^{\prime}\rightarrow H^{\prime}$ defined by $T:=\iota\circ\widetilde{A}^{-1}$. This operator is compact, self-adjoint and positive because
$(T\mathbf{f},\mathbf{g})_{H^{\prime}}=\langle\iota(\widetilde{A}^{-1}\mathbf{f}),\widetilde{A}^{-1}\mathbf{g}\rangle_{H^{\prime}\times H}=\int_{\Gamma}\widetilde{A}^{-1}\mathbf{f}\cdot\widetilde{A}^{-1}\mathbf{g}=(\mathbf{f},T\mathbf{g})_{H^{\prime}},$
and $(T\mathbf{f},\mathbf{f})_{H^{\prime}}=\|\mathbf{f}\|_{H^{\prime}}^{2}\geq 0$ for each $\mathbf{f},\mathbf{g}\in H^{\prime}$.
We are now ready to establish the spectral properties of $\widetilde{A}:H\rightarrow H^{\prime}$.
Theorem 4.10 (eigendecomposition of the Stokes operator).
Assume $\Gamma$ is of class $C^{2}$. Then, there exists a family of functions $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}\subseteq\bigcap_{2\leq p<\infty}\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ and a nondecreasing sequence of $\{\omega_{n}\}_{n\in\mathrm{N}}$ of positive numbers with $\lim_{n\to\infty}\omega_{n}=\infty$ such that
  1. (i)
    For each $n\in\mathrm{N}$, $(\mathbf{v}_{n},\omega_{n})$ is an eigenpair of the Stokes operator, i.e.: $\widetilde{A}\mathbf{v}_{n}=\omega_{n}\iota(\mathbf{v}_{n})$ in $(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$, or equivalently,
    $\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}_{n}:\nabla_{\Gamma}\mathbf{w}=\omega_{n}\int_{\Gamma}\mathbf{v}_{n}\cdot\mathbf{w},\qquad\forall\mathbf{w}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma).$ (4.21)
  2. (ii)
    $\{\omega_{n}^{1/2}\iota(\mathbf{v}_{n})\}_{n\in\mathrm{N}}$ is an orthonormal basis of $H^{\prime}=(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$ with respect to the equivalent inner product (4.20).
  3. (iii)
  4. (iv)
    $\{\omega_{n}^{-1/2}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is an orthonormal basis of $H=\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ with respect to the equivalent inner product (4.19).
Proof.
We already know that the linear operator $T$ is compact, self-adjoint and positive with respect to the inner product (4.20). Hence, since $(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$ is separable, it follows from [27, §D.6] that there exist an orthogonal basis $\{\mathbf{z}_{n}\}_{n\in\mathrm{N}}$ of $(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$ and a sequence of (positive) nonincreasing numbers $\{\lambda_{n}\}_{n\in\mathrm{N}}$ with $\lim_{n\to\infty}\lambda_{n}=0$ such that $T\mathbf{z}_{n}=\lambda_{n}\mathbf{z}_{n}$ in $(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$. Equivalently, upon defining $\mathbf{v}_{n}:=\widetilde{A}^{-1}\mathbf{z}_{n}\in\mathbf{H}^{1}_{t,\sigma}(\Gamma)$ and $\omega_{n}:=\lambda_{n}^{-1}$, we see that $\widetilde{A}\mathbf{v}_{n}=\omega_{n}\iota(\mathbf{v}_{n})$ in $(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}$ as well as
@@ 4.4. The Stokes operator
which proves (4.21).
Without loss of generality assume $(\mathbf{z}_{n},\mathbf{z}_{k})_{(\mathbf{H}^{1}_{t,\sigma}(\Gamma))^{\prime}}=\delta_{nk}\omega_{k}$ for each $n,k\in\mathrm{N}$. To show that $\{\omega_{n}^{1/2}\mathbf{v}_{n}\}_{n\in\mathrm{N}}\equiv\{\omega_{n}^{1/2}\iota\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is orthonormal in $H^{\prime}$, it is enough to notice that
$(\iota\mathbf{v}_{n},\iota\mathbf{v}_{k})_{H^{\prime}}=(\iota\widetilde{A}^{-1}\mathbf{z}_{n},\iota\widetilde{A}^{-1}\mathbf{z}_{k})_{H^{\prime}}=(T\mathbf{z}_{n},T\mathbf{z}_{k})_{H^{\prime}}=\omega_{n}^{-1}\omega_{k}^{-1}(\mathbf{z}_{n},\mathbf{z}_{k})_{H^{\prime}}=\omega_{n}^{-1}\delta_{nk},$
for each $n,k\in\mathrm{N}$. In turn, for the $H$-orthonormality of $\{\omega_{n}^{-1/2}\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ and the $\mathbf{L}^{2}_{t,\sigma}(\Gamma)$-orthonormality of $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ it suffices to see that
[…]
To show that $\{\iota\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is a basis of $H^{\prime}$, let $\mathbf{f}\in H^{\prime}$ such that $(\mathbf{f},\iota\mathbf{v}_{n})_{H^{\prime}}=0$ for each $n\in\mathrm{N}$. That is, $(\mathbf{f},T\mathbf{z}_{n})_{H^{\prime}}=\lambda_{n}(\mathbf{f},\mathbf{z}_{n})_{H^{\prime}}=0$ for each $n\in\mathrm{N}$. Since $\lambda_{n}\neq 0$ and $\{\mathbf{z}_{n}\}_{n\in\mathrm{N}}$ is an orthogonal basis of $H^{\prime}$, we deduce that $\mathbf{f}=\mathbf{0}$. Hence, $\operatorname{span}\{\omega_{n}^{1/2}\iota\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ is dense in $H^{\prime}$. Items () iii and () iv are shown in a similar fashion, and hence we omit their proofs.
Finally, for each $p\in(2,\infty)$, the $\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$-regularity of each $\mathbf{v}_{n}$ follows from a bootstrapping argument resulting from combining the invertibility of the Stokes operator (4.18) applied to problem (4.22) with Sobolev embeddings (cf. Section 2.1). ∎
As expected, if $\Gamma$ is more regular, say of class $C^{2,1}$, then the family $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}$ provided by 4.10 are actually a.e. eigenfunctions of the “strong” Stokes operator $A:=-P_{\sigma}(\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\cdot):\mathbf{W}^{2,p}_{t,\sigma}(\Gamma)\rightarrow\mathbf{L}^{p}_{t,\sigma}(\Gamma)$; see 4.11 below. Before stating the result notice that integration-by-parts formula (2.5) and the definition of the Leray projection (cf. 4.4) let us write
$\langle\widetilde{A}\mathbf{u},\mathbf{v}\rangle_{(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}\times\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}\cdot\nabla_{\Gamma}\mathbf{v}=\int_{\Gamma}A\mathbf{u}\cdot\mathbf{v},\qquad\forall\mathbf{u}\in\mathbf{W}^{2,p}_{t,\sigma}(\Gamma),\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma),$ (4.24)
Consequently, if $\Gamma$ is of class $C^{m+2,1}$ for some $m\in\mathrm{N}\cup\{0\}$, for each $\mathbf{u}\in\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$ and $\mathbf{f}\in\mathbf{W}^{m,p}_{t}(\Gamma)$, it follows that: $\widetilde{A}\mathbf{u}=P_{\sigma}\mathbf{f}$ in $(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ if and only if $A\mathbf{u}=P_{\sigma}\mathbf{f}$ a.e. on $\Gamma$, or equivalently $P_{\sigma}(\mathbf{f}+\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}\mathbf{u})=\mathbf{0}$. Invoking 4.4 (Leray projection in $\mathbf{L}^{p}_{t}(\Gamma)$) along with (4.6), there exists a unique $\uppi\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ such that
@@
Moreover, if $\mathbf{f}\in\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$, then by 4.8 (higher $\mathrm{L}^{p}$-based regularity of the tangent Stokes equations) we deduce that $A$ is a continuous bijection from $\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$ to $\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)$ and
$\|\mathbf{u}\|_{m+2,p;\Gamma}+\|\uppi\|_{m+1,p;\Gamma}\lesssim\|\mathbf{f}\|_{m,p;\Gamma}=\|A\mathbf{u}\|_{m,p;\Gamma}.$ (4.25)
Conversely, the Bounded Inverse Theorem [12, Corollary 2.7] (see also [30, Teorema 3.4]) implies that
$\|A\mathbf{u}\|_{m,p;\Gamma}\lesssim\|\mathbf{u}\|_{m+2,p;\Gamma}.$ (4.26)
@@
defines an equivalent inner product in $\mathbf{H}^{m+2}_{t,\sigma}(\Gamma)$. On the other hand, if $m+2=2\ell+1$ is odd we also use 3.3 (Poincaré inequality) to deduce that
@@
defines an equivalent inner product in $\mathbf{H}^{m+2}_{t,\sigma}(\Gamma)$; we refer to [12, Section 7.3] for similar constructions. The following result shows that, provided $\Gamma$ is sufficiently regular, the eigenpairs $\{(\mathbf{v}_{n},\omega_{n})\}_{n\in\mathrm{N}}$ provided by 4.10 enjoy higher regularity and orthogonality with respect to the inner products (4.27) and (4.28).
Theorem 4.11 (higher regularity of the eigenfunctions of the Stokes operator).
Assume $\Gamma$ is of class $C^{m+2,1}$ for some nonnegative integer $m$. Then, the family of eigenpairs $\{(\mathbf{v}_{n},\omega_{n})\}_{n\in\mathrm{N}}$ provided by 4.10 satisfy
@@
    $\{\mathbf{v}_{n}\}_{n\in\mathrm{N}}\subseteq\bigcap_{2\leq p<\infty}\mathbf{W}^{m+2,p}_{t,\sigma}(\Gamma)$.